payton:

1/2 (x-14) + 11= 1/2 x - (x - 4) What is the value of x ? –1 –1/2 0 1/2

justus:

\[\frac{ 1 }{ 2 }(x-14)+11=\frac{ 1 }{ 2}x-x\] \[\frac{ x-14 }{ 2}+11=\frac{ 1 }{ 2 }x-x+4\] \[\frac{ x-14 }{ 2}+11=\frac{ x }{ 2 }-x+4\] \[\frac{ x-14 }{ 2}+11=4-\frac{ x }{ 2 }\] \[x-14+22=8-x\] \[x+8=8-x\] \[x=-x\] \[x+x=0\] \[2x=0\] Divide both sides by 2 and get your answer :)